OpenStudy (anonymous):

solve the system of equations using the graphing method {4x-2y=-4 {3x+y=7

OpenStudy (anonymous):

what type of math class

OpenStudy (anonymous):

4x-2y=-4 3x+y=7 ______________ 4x-2y=-4 6x+2y=14 ____________ 10x=10 --------- @chrystaldenise :Got it ?:) If got find y :)

OpenStudy (anonymous):

x=1

OpenStudy (anonymous):

nope still confuse

OpenStudy (anonymous):

First u should to know u want to find x or y ? What do u want to find first ?:)

OpenStudy (anonymous):

i guess y

OpenStudy (anonymous):

K ! Now to find it we should kill the X s !

OpenStudy (anonymous):

for the first line i did4x-2y=-4 -4x -4x -2y=-4x-4 /-2= /-2 y=-4x/-2 = 2 -4/-2 = 4

OpenStudy (anonymous):

k ?

OpenStudy (anonymous):

ok

OpenStudy (xishem):

"using the graphing method" This isn't using the graphing method. Graph the two functions, and the points where they cross each other are solutions to the system of equations.

OpenStudy (anonymous):

so solve them both and than point the solutions